### v2

# Learning to Intervene on Concept Bottlenecks

Abstract
While traditional deep learning models often lack interpretability, concept bottleneck models (CBMs) provide inherent explanations via their concept representations. Specifically, they allow users to perform interventional interactions on these concepts by updating the concept values and thus correcting the predictive output of the model. Traditionally, however, these interventions are applied to the model only once and discarded afterward. To rectify this, we present concept bottleneck memory models (CB2M), an extension to CBMs. Specifically, a CB2M learns to generalize interventions to appropriate novel situations via a two-fold memory with which it can learn to detect mistakes and to reapply previous interventions. In this way, a CB2M learns to automatically improve model performance from a few initially obtained interventions. If no prior human interventions are available, a CB2M can detect potential mistakes of the CBM bottleneck and request targeted interventions. In our experimental evaluations on challenging scenarios like handling distribution shifts and confounded training data, we illustrate that CB2M are able to successfully generalize interventions to unseen data and can indeed identify wrongly inferred concepts. Overall, our results show that CB2M is a great tool for users to provide interactive feedback on CBMs, e.g., by guiding a user’s interaction and requiring fewer interventions.
1 Introduction
Deep learning models often represent black-box models that make it difficult for human users to understand their decision processes Adadi & Berrada (2018); Cambria et al. (2023); Saeed & Omlin (2023) and interact with these Schramowski et al. (2020); Teso et al. (2023). One recent branch within eXplainable Artificial Intelligence (XAI) focuses on the potential of so-called concept bottleneck models (CBMs) Koh et al. (2020); Stammer et al. (2021) to tackle such issues.
These are designed to be inherently human-interpretable. They perform inference (e.g., for bird image classification, cf. Fig. 1 (top)), by transforming the initial raw input into a set of human-understandable concepts (e.g., wing shape and color) with a bottleneck network and provide a final task prediction based on the activation of these concepts with a predictor network. The concept activations thereby serve as an inherent explanation of the model’s decision Teso et al. (2023). Arguably even more valuable, these concept activations can be used as a means for humans to perform targeted interactions, e.g., for querying further explanations Abid et al. (2022) or to correct the model’s concept prediction Koh et al. (2020).
A recent surge of research has focused on the benefits of leveraging interactions in AI models in general Ouyang et al. (2022); Miller (2019), and also CBMs in particular Teso et al. (2023). Multiple such approaches focus on leveraging interactions for mitigating errors of the predictor network Bontempelli et al. (2021); Stammer et al. (2021). However, little work has focused on mitigating potential errors of the initial bottleneck network. Moreover, although interventional interactions on a CBM’s concept activations are a natural tool for this purpose, they have received little attention since their introduction by Koh et al. Koh et al. (2020).
One likely reason for this is that interventions according to Koh et al. (2020) represent a once-use tool for updating model performance by adding human-provided concept labels to an increasing number of randomly selected concepts. For sustainably improving a model’s performance, however, this approach is inefficient and potentially demands a large number of repeated user interactions, where providing such repeated feedback has been identified to lead to a loss in focus of human users Amershi et al. (2014).
In this work, we therefore argue to harvest the rich information present in previously collected interventions in a multi-use approach. Specifically, let us suppose a user corrects a model’s inferred concepts through a targeted intervention. In that case, the intervention carries information on where the model did not perform well. This can be used to improve prediction in similar future situations (Fig. 1 bottom). In this context, we therefore introduce Concept Bottleneck Memory Models (CB2M) as a novel and model-agnostic extension to CBMs. CB2M are based on adding a two-fold memory of interventions to the CBM architecture, which allows to keep track of previous model mistakes as well as previously applied interventions. This memory enables a CB2M to reapply interventions when the CBM repeats mistakes and thus automatically corrects them without the need for additional human feedback. This ultimately allows to overcome the issue of one-time interventions of standard CBMs and enables the model to learn more effectively from provided human feedback. Overall, however, human feedback can be unavailable, and obtaining it is costly. CB2M mitigates this issue by its ability to detect potential model mistakes prior to initial human feedback where its memory module can be used to specifically select data points for human inspection, and thus guide human feedback to where it is really needed.
Figure 1: Reusing a CBM intervention can correct model mistakes for multiple examples. Top: CBMs generate a human interpretable concept representation to solve the final task. Human users can correct these concept predictions via targeted interventions (blue) influencing the final prediction. Bottom: Human interventions hold valuable information that can be reused in suitable situations to automatically correct model errors without further human interactions.
We illustrate the full potential of CB2M in our experimental evaluations on several challenging tasks, such as handling distribution shifts and confounding factors across several datasets. In summary, we make the following contributions: (i) We identify the potential of extracting generalizable knowledge from human interventions as a means of correcting concept bottleneck models. (ii) We introduce CB2M, a flexible extension to any CBM-like architecture for handling such interactive interventions. (iii) Our experimental evaluations show that CB2M can truly learn from interventions by generalizing them to previously unseen examples. (iv) We further show that CB2M are also able to detect model mistakes without the need for initial human knowledge and thus allow to query a user for targeted interventions.
The rest of the paper proceeds as follows: In Sec. 2, we provide a brief background followed by the introduction of CB2M. We present our experiments evaluations in Sec. 3. Afterwards, we relate CB2M to other work in Sec. 4 before concluding the paper together with potential future research directions in Sec. 5.
2 Concept Bottleneck Memory Models (CB2M)
Let us first introduce the background notations on CBMs and interventions before introducing CB2M for detecting model mistakes and generalizing interventions to novel, unseen examples.
2.1 Background
A CBM which solves the task of transforming inputs $\mathcal{X}$ to outputs $\mathcal{Y}$ consists of two parts. The bottleneck model $g:x\rightarrow c$ transforms an input $x\in\mathcal{X}$ into its concept representation $c$. Afterward, the predictor network $f:c\rightarrow y$ uses this representation to generate the final target output $y\in\mathcal{Y}$. The correct, i.e., ground-truth values for $c$ and $y$ are written as $c^{*}$ and $y^{*}$, respectively. We refer to overall model (task) accuracy as $\text{Acc}_{f}$ and to concept accuracy as $\text{Acc}_{g}$.
Human interactions with the concept representations are called interventions. An intervention $i\in\mathcal{I}$ is a set of tuples $i=\{(c^{\prime}_{j},j)|j\in\mathcal{J}_{i}\}$, with updated concept values $c^{\prime}_{j}$ and concept indices $j$. $\mathcal{J}_{i}$ is the set of all indices for intervention $i$. Applying an intervention to a sample $x$ overwrites the predicted concept values with those of the intervention, which we denote as $x|i$.
As CBMs consist of two processing modules, the bottleneck and predictor networks, errors can occur in either, with different consequences on how to handle these Bontempelli et al. (2021). If the bottleneck makes an error, this error will most likely also negatively influence the predictor. On the other hand, it is also possible that the predictor makes a wrong final prediction despite having received a correct concept representation. In the latter case, the concept space is either insufficient to solve the task, or the predictor network is susceptible to, e.g., some spurious correlations. Where other works have investigated handling an insufficient concept space through additional (unsupervised) concepts Sawada & Nakamura (2022), or correcting a predictor with spurious correlations Stammer et al. (2021) CB2M focuses on mitigating errors that originate from the bottleneck model. This is achieved by utilizing interventions on the concept space. Let us discuss this in more detail in the following.
2.2 Concept Bottleneck Memory Models
Figure 2: Overview of CB2M to detect mistakes or generalize interventions. A vanilla CBM (grey) is extended with a two-fold memory (orange and green). The memory compares encodings of new samples to known mistakes to (i) detect model errors or (ii) automatically correct the model via reuse of interventions.
Let us now introduce Concept Bottleneck Memory Models (CB2M) as a flexible extension for any CBM architecture. The bottleneck and predictor networks of the CBM remain unchanged but are extended by a two-fold memory module $\mathcal{M}$ which consists of a mistake memory $\mathcal{M}^{m}$ coupled with a intervention memory $\mathcal{M}^{i}$. The mistake memory operates on encodings $x_{e}$, i.e., the input of the last layer of the bottleneck network. It measures the similarity between two data points $x$ and $x^{\prime}$ via the euclidean distance of their encodings, $d(x_{e},x^{\prime}_{e})=\|x_{e}-x^{\prime}_{e}\|$. The intervention memory directly keeps track of known interventions and associates them to elements of the mistake memory, meaning that the memorized intervention $i$ can be used to correct the memorized mistake of $x_{e}$. We denote an associated encoding and intervention as $\alpha(x_{e},i)$. Overall, this joint memory can be used to detect model mistakes (orange in Fig. 2) or enable automatic reuse of memorized interventions (green in Fig. 2), which we explain in detail in the following paragraphs.
By extending the vanilla CBM with a memory, CB2M can be used for two distinct tasks (cf. Fig. 2): (i) detecting potential model mistakes and (ii) generalizing interventions to new examples. Besides the general advantage of knowing when an AI model has made an incorrect prediction, this knowledge is even more relevant for CBMs as this can be utilized to focus human attention toward obtaining a targeted intervention. Thus, with the ability to handle task (i) CB2M is especially relevant when humans want to provide interventional feedback to a CBM. Furthermore, after humans have intervened on a CBM, they have, in fact, provided valuable knowledge also for future situations. Thus, an applied intervention carries information on how to correct a particular model mistake. We claim that this information should not be discarded as in the original work of Koh et al. Koh et al. (2020), but be reused when similar mistakes occur again. This is where task (ii) of CB2M comes into play.
Detecting Wrongly Classified Instances.
Intuitively, if a data point is similar to other examples where the model has made mistakes, the model will more likely repeat these mistakes on the new data point. Therefore, we utilize the mistake memory $M_{m}$ to keep track of previous mistakes (cf. Alg. 1 for pseudo-code). First, the memory is filled with encodings of datapoints, for which the model did not initially generate the correct output and for which the concept accuracy is smaller than a threshold $t_{a}\in[0,1]$. This leads to: $\mathcal{M}^{m}=\{x_{e}:f(g(x))\neq y^{*}\wedge Acc_{g}(x)<t_{a}\}$. For a new unseen instance $\hat{x}$, we then compare its encoding $\hat{x}_{e}$ with the mistakes in the memory $\mathcal{M}^{m}$ (Alg. 1, lines 4 - 8). If we find $k$ mistakes with a distance to $\hat{x}_{e}$ smaller than $t_{d}$, we consider a model to be making a known mistake. Formally, we thus predict a model mistake for a new unseen instance $\hat{x}$ if:
$\forall j\in\{1,~{}...~{},k\}:\exists x_{e,j}\in\mathcal{M}^{m}:d(\hat{x}_{e},x_{e,j})\leq t_{d}$ (1)
This mistake memory can initially be filled with known model mistakes, for example, based on the validation set. However, once the CB2M is in use, the memory of mistakes will continuously be updated via interactive feedback, and new encodings will be added. This can constantly improve detection during deployment as corrective interventions can immediately be requested after detecting a potentially misclassified sample.
1:  Memory setup: $\mathcal{M}^{m}\leftarrow\{x_{e}:x\in\mathcal{X}_{val}\wedge f(g(x))\neq y^{*}\wedge Acc_{g}(x)<t_{a}\}$
2:  $\hat{x}\leftarrow\texttt{New unseen instance};$ $j\leftarrow 0$
3:  for $m\in\mathcal{M}^{m}$ do
4:     if $d(\hat{x}_{e},m)\leq t_{d}$ then
5:        $j\leftarrow j+1$
6:     end if
7:  end for
8:  if $j\geq k$ then
9:     return Mistake
10:  else
11:     return No mistake
12:  end if
Algorithm 1 Detection of Model Mistakes. Given: Parameters $t_{d}$, $t_{a}$ and $k$, data set for memory setup (e.g. validation set) $\mathcal{X}_{val}$ and a CBM with bottleneck $f$ and predictor $g$.
Generalization of Interventions.
Next to detecting model errors with the mistake memory, we can use both the mistake memory and the intervention memory together to generalize interventions. As initially introduced in Koh et al. (2020), interventions for correcting predicted concept activations only apply to a single sample. However, we claim that these interventions also contain valuable information for further samples and should thus be reused, thereby reducing the need for additional future human interactions. Intuitively, if an intervention is applicable for one example, it is likely also relevant for similar inputs, at least to a certain degree.
To achieve such intervention generalization from one sample to several, we utilize both parts of the CB2M memory. Specifically, whenever an intervention $i$ is applied to a model, we store it in the intervention memory $\mathcal{M}^{i}$ and keep the encoding of the original input point in the mistake memory $\mathcal{M}^{m}$. We also keep track of the corresponding entries $\alpha(x_{e},i)$. When the model receives a new sample $\hat{x}$, we next check for similar encodings in the mistake memory $\mathcal{M}^{m}$ according to Eq. 1. Here, we use $k=1$ and only consider the most similar mistake and its interventions. If there is indeed a mistake encoding $x_{e}$ within distance $t_{d}$ of $\hat{x}_{e}$, we apply its associated intervention $i$ (with $\alpha(x_{e},i)$) to the new data point $\hat{x}$. If there is no similar mistake, we let the model perform its prediction as usual.
The threshold $t_{d}$ is crucial for intervention generalization, as it directly controls the necessary similarity to reapply memorized interventions. Selecting a suitable value for $t_{d}$ differs from the mistake prediction use case as we want to generalize as many interventions as possible under the constraint that the generalized interventions remain valid. To this end, we call an intervention $i$ for a sample $x$ valid if the class prediction after intervening is not worse than before. We write this as $val(x,i):f(g(x))=y^{*}\implies f(g(x|i))=y^{*}$. Given that, we want to maximize $t_{d}$, while the following remains true:
$\forall x,x^{\prime}\in\mathcal{X}:d(x_{e},x_{e}^{\prime})\leq t_{d}\Rightarrow\forall i\in\mathcal{I}:val(x,i)\Rightarrow val(x^{\prime},i)$ (2)
We can also express this in terms of full dataset accuracy, where our dataset accuracy after applying interventions should be larger than (or equal to) the accuracy without interventions:
$Acc_{f}(\mathcal{X}|\mathcal{M})\geq Acc_{f}(\mathcal{X}).$ (3)
Here $\mathcal{X}|\mathcal{M}$ is the dataset $\mathcal{X}$ with applied interventions from the memory $\mathcal{M}$:
$\begin{split}\mathcal{X}|\mathcal{M}=&\{x|i:x\in\mathcal{X}:\exists x_{e}^{\prime}\in\mathcal{M}^{m}:\exists i\in\mathcal{M}^{i}:d(x_{e},x^{\prime}_{e})\leq t_{d}\wedge\alpha(x_{e}^{\prime},i)\}\\ &\cup\{x:x\in\mathcal{X}:\neg\exists x_{e}^{\prime}\in\mathcal{M}^{m}:d(x_{e},x^{\prime}_{e})\leq t_{d}\}\end{split}$ (4)
Thus, we want to find the largest $t_{d}$ that still satisfies Eq. 3. To do that, we can set up the memory $\mathcal{M}$ based on the validation set by adding all model mistakes to $\mathcal{M}^{m}$ and simulating corresponding interventions with ground-truth labels for $\mathcal{M}^{i}$. The selection of $t_{d}$ is then done on the training set. This results in $\mathcal{M}^{m}=\{x_{e}:x\in\mathcal{X}_{val}\wedge f(g(x))\neq y^{*}\}$ and $\mathcal{M}^{i}=\{i:i\in\mathcal{I}\wedge x_{e}\in\mathcal{M}^{m}\wedge\alpha(x_{e},i)\wedge\forall j\in\mathcal{J}_{i}:c^{\prime}_{j}=c^{*}_{j}\}$.
2.3 General Perspective
Assumptions About Human Feedback.
A CB2M leverages human feedback to improve upon the CBM via interventions. To this end, it is assumed that the feedback provided by humans is correct. This is a common assumption in work on CBMs Koh et al. (2020); Chauhan et al. (2022) and (inter)active learning in general Settles (2009); Berg et al. (2019). As this assumption does not always hold true, we delve deeper into it in the focus of CBMs. Often, it is easier for humans to provide concept information than to provide information on the complete task. For example, when considering bird species classification cf. Fig. 1, it is easier to identify the bird’s color than its species. This phenomenon occurs when concepts are "low-level" and human-understandable. In other domains, such as the medical one, providing correct concept labels may require expert domain knowledge, but it is still possible and easier to infer concept labels than class labels. Disregarding a human’s ability to provide correct feedback, e.g., concept labels, this does not imply they will. A user with malicious intentions could actively provide wrong concept labels to misguide the system. This potential danger has to be considered when incorporating human feedback, i.e., also in the context of CB2M. Recent work has begun tackling this issue Ju et al. (2022).
Handling Bottleneck Errors via CB2M.
As stated above, while CBMs have many advantageous properties compared to standard deep models, their structure can also make error analysis more difficult due to separate processing via the bottleneck and predictor networks Marconato et al. (2023). CB2M stands in line with research that tackles error sources in CBMs separately Bontempelli et al. (2021). Where several previous works have tackled mitigating errors in the predictor network Sawada & Nakamura (2022); Stammer et al. (2021); Teso et al. (2023), CB2M is designed to tackle errors of the bottleneck network via interventions. These approaches, overall, stand in contrast to techniques that consider the errors of the whole model Jiang et al. (2018); Hendrycks & Gimpel (2017). However, our evaluations particularly support the claim of Bontempelli et al. Bontempelli et al. (2021) on the advantages of dividing the treatment of different error sources in CBMs. As such, CB2M should be seen as an important tool for this goal, where other methods are required for specific errors of the predictor network.
3 Experimental Evaluations
To evaluate the potential of CB2M in intervention generalization and mistake detection, we provide evidence via several evaluations. First, we evaluate the ability of CB2M to detect the existence of similar data points. Afterward, we continue to more challenging scenarios and investigate CB2M in the context of unbalanced and confounded data as well as data affected by distribution shifts. Let us first describe the experimental setup.
3.1 Setup
Data: The Caltech-UCSD Birds (CUB) dataset Wah et al. (2011) consists of 11788 images of 200 different bird classes. We use the data splits provided by Koh et al. (2020), resulting in training, validation, and test sets with 40, 10, and 50% of the total images. Additionally, we add 4 more folds of training and validation set to perform 5-fold validation. Images in the dataset are annotated with 312 concepts (e.g., beak-color:black, beak-color:brown, etc.), which can be grouped into concept groups (one group for all beak-color:_ concepts). We follow the approach of previous work Koh et al. (2020); Chauhan et al. (2022) and use only concepts that occur for at least 10 classes and then perform majority voting on the concept values for each class. This results in 112 concepts from 28 groups.
We further provide evidence based on the MNIST LeCun & Cortes (1998), confounded ColorMNIST (C-MNIST) Rieger et al. (2020) and SVHN Netzer et al. (2011) datasets. For all three, we train the model for the parity MNIST task as in Mahinpei et al. (2021). Hereby, the digit in the image is considered the concept, and the class label describes whether the digit is even or odd. Furthermore, rather than evaluating on the original MNIST dataset, we focus on an unbalanced version of this task. In this setting, we remove 95% of the training data of one class (for the results in the main paper, the digit “9”, for other digits cf. App. A.4). We refer to App. A.3 for results on the original MNIST dataset, indicating that current base models yield very high performances and make additional interventions unnecessary. Lastly, for C-MNIST, each digit is colored in the same color during model training (the confounder), but the coloring is random during test time.
We use the standard train and test splits for these datasets and create validation sets with 20% of the training data. As for CUB, we generate 5 training and validation folds in total. When considering human interventions, we follow the common assumption that humans provide the correct concept values as long as the requested concepts are present in the input (e.g., visible in an image).
Models: For CUB, we use the same model setup as Koh et al. (2020), instantiating the bottleneck model with the Inception-v3 architecture Szegedy et al. (2016) and the predictor network with a simple multi-layer perceptron. For MNIST variants and SVHN, we use MLPs for bottleneck and predictor networks as in Mahinpei et al. (2021). All CBMs are trained with the independent scheme, as this has shown to be the most effective model training regarding interventions Shin et al. (2023). Further training details can be found in App. A.1. We use CB2M as described in Sec. 2.2 to enable the generalization of interventions and detection of model mistakes. CB2M parameters are tuned for generalization and detection separately on the training and validation set. For all detection experiments, the memory of CB2M is filled with wrongly classified instances of the validation set according to the parameters. For generalization experiments, we simulate human interventions on the validation set and use CB2M to generalize them to the test set. Whenever we consider interventions on a subset of concepts or concept orderings, we use ECTP Shin et al. (2023).
Metrics: We evaluate CB2M and the base model with several different metrics. We use both concept and class accuracy of the underlying CBM (with and without CB2M) to observe improvements in the final task and to investigate the intermediate concept representation. We evaluate the detection of model mistakes using AUROC and AUPR, in line with related work Ramalho & Miranda (2019). To observe how interventions improve model performance, we propose normalized relative improvement (NRI), which measures improvement independent of different baseline values. NRI measures the percentage of the maximum possible improvement in class accuracy achieved:
$\text{NRI}=\frac{\Delta}{\Delta_{\text{max}}}=\frac{\text{Acc}_{f}-\text{Acc}_{f,\text{base}}}{\text{Acc}_{f,\text{max}}-\text{Acc}_{f,\text{base}}}$ (5)
Where $\text{Acc}_{f}$ ($\text{Acc}_{f,\text{base}}$) refers to the model accuracy after (before) applying interventions and $\text{Acc}_{f,\text{max}}$ is the maximum possible accuracy to achieve through interventions. To apply this metric, $\text{Acc}_{f,\text{max}}$ has to be known. To estimate this, one can use, e.g., the accuracy of the predictor network given ground-truth concept information on the validation set.
Table 1: CB2M generalizes interventions to unseen data points. Performance of CBM and CB2M on data points identified for generalized interventions and on the full datasets. Over all datasets, generalizing interventions with CB2M improves concept accuracy substantially. This further results in a drastic improvement of class accuracy on identified data points and a noticeable improvement on the full set. (Best values bold, average and standard deviation over augmented test set versions (CUB) or 5 runs (other)).
Concept Acc. ($\uparrow$) Class Acc. ($\uparrow$)
Dataset Setting CBM CB2M CBM CB2M
CUB Identified $86.4\pm 2.7$ $\mathbf{99.0}\pm 0.7$   $5.0\pm 1.7$ $\mathbf{88.7}\pm 5.4$
Full set $94.7\pm 0.6$ $\mathbf{98.7}\pm 3.5$ $64.8\pm 2.7$ $\mathbf{69.1}\pm 5.5$
Parity MNIST Identified $85.3\pm 2.6$ $\mathbf{98.7}\pm 0.4$ $22.5\pm 5.7$ $\mathbf{93.7}\pm 1.9$
(unbalanced) Full set $97.5\pm 0.2$ $\mathbf{98.0}\pm 0.3$ $91.2\pm 0.1$ $\mathbf{94.0}\pm 1.2$
Parity C-MNIST Identified $82.2\pm 0.6$ $\mathbf{95.5}\pm 1.2$ $20.1\pm 7.1$ $\mathbf{85.9}\pm 4.7$
Full set $87.1\pm 0.0$ $\mathbf{88.4}\pm 0.4$ $68.6\pm 0.3$ $\mathbf{74.9}\pm 2.1$
3.2 Results
Beyond One-Time Interventions.
First, we analyze how well CB2M generalizes interventions to unseen data points. If a standard CBM receives a new input similar to a previous datapoint with a corresponding intervention, that intervention is not used at all. CB2M, on the other hand, allows the reuse of information provided in previous interventions. To evaluate the generalization in these cases, we provide results on the CUB dataset where, due to its small number of images per class, we expand the dataset by applying the following data augmentations to the images: color jitter, blurring, blackout, as well as salt&pepper, and speckles noise. In this way, these augmented images act as a proxy for real-world similar inputs, e.g., images with different lighting conditions. We thus fill CB2M with simulated human interventions on the test set and evaluate how they generalize to a novel augmented test set version. Tab. 1 shows that CB2M substantially improves upon the base CBM, both on instances identified for intervention generalization and on the full set.
Next, we evaluate CB2M under more challenging settings, namely training with highly unbalanced or confounded training data. As seen in Tab. 1 the base CBM struggles to learn the underrepresented digit in the unbalanced Parity MNIST dataset. On the confounded Parity C-MNIST dataset [^1] For this dataset, we assume that we have access to some human interventions on unconfounded data. we observe that the standard CBM is strongly influenced by the confounding factor which negatively impacts the bottleneck performance during test time. By generalizing from few human interventions CB2Ms, on the other hand, can substantially improve performance compared to the vanilla CBM. In both cases, the reapplied interventions reach a concept accuracy close to 100%, showing that the interventions successfully correct the bottleneck errors. Correcting the concept representation on the instances identified for reapplied intervention also substantially boosts the class accuracy on these instances. Overall, we can conclude that CB2M is able to generalize intervention successfully. This holds true for naturally similar inputs and scenarios like unbalanced and confounded data.
Asking for Interventions.
Next, we go from the generalization of interventions to the second use-case for which CB2M can be deployed, namely for detecting model mistakes prior to human feedback. For this, we compare CB2M to two baselines. The random baseline for mistake detection simply marks random samples as mistakes. In contrast, softmax based detection of mistakes uses the softmax probability of the strongest activated class as a proxy to predict whether the model made a mistake Hendrycks & Gimpel (2017). Where the softmax baseline uses information from the end of the model, i.e., after the predictor network, CB2Ms estimate model errors only based on the bottleneck network. While detecting mistakes of the whole model covers all potential model errors, we hypothesize that detecting mistakes of the bottleneck network is more suitable for interventions, as they are tied to the bottleneck network. We compare CB2M to the baselines on CUB and the Parity MNIST (unbalanced) datasets. Additionally, we evaluate the detection on Parity C-MNIST, where the methods have access to a small number of unconfounded data points.
Table 2: CB2M detects wrongly classified instances. AUROC and AUPR values on the test set. For the confounded Parity C-MNIST, CB2M can even achieve substantially better detection than the baselines. (Best values bold, average and standard deviations over 5 runs.)
Dataset Confounded Metric Random Softmax CB2M
CUB No AUROC ($\uparrow$) $51.1\pm 0.7$ $83.7\pm 1.1$ $\mathbf{84.8}\pm 0.7$
AUPR ($\uparrow$) $77.3\pm 0.4$ $94.0\pm 0.6$ $\mathbf{94.6}\pm 0.3$
Parity MNIST No AUROC ($\uparrow$) $50.5\pm 0.1$ $\mathbf{90.7}\pm 1.7$ $88.7\pm 0.4$
(unbalanced) AUPR ($\uparrow$) $91.2\pm 0.1$ $\mathbf{98.8}\pm 0.3$ $98.5\pm 0.1$
Parity C-MNIST Yes AUROC ($\uparrow$) $50.3\pm 0.7$ $65.7\pm 0.3$ $\mathbf{83.4}\pm 0.8$
AUPR ($\uparrow$) $69.0\pm 0.6$ $79.8\pm 0.3$ $\mathbf{91.5}\pm 0.4$
Table 3: Interventions based on CB2M detection successfully improve model performance. NRI of interventions on identified instances and the full test set. Samples for intervention identified by the three methods. As expected, interventions improve performance on selected instances for all three methods. More importantly, using CB2M leads to considerably larger improvements on the full dataset. (Best values bold, standard deviations over 5 runs.) Setting Random Softmax CB2M CUB Identified    $95.4\pm 0.6$    $\mathbf{96.3}\pm 0.6$    $95.9\pm 0.5$ Full Set    $34.3\pm 5.7$    $70.1\pm 3.1$    $\mathbf{75.5}\pm 4.5$ Parity MNIST (unbalanced) Identified $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ Full Set    $13.2\pm 4.2$    $62.1\pm 4.9$    $\mathbf{69.6}\pm 4.1$ Parity C-MNIST Identified $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ Full Set    $60.0\pm 9.8$    $87.3\pm 0.8$    $\mathbf{89.7}\pm 6.1$
Figure 3: Less is more: Intervening on a subset of all concepts already yields large improvements. Interventions on identified instances by CB2M on CUB achieve large portions of the total improvement by interventions on less than 50% of concept groups. Average and standard deviation over 5 runs, concept ordering computed with ECTP.
First, in Tab. 2, we show that the mistake detection of CB2M performs on par with the softmax baseline for CUB and Parity MNIST (unbalanced) while even outperforming it for Parity C-MNIST. On the latter, CB2M is able to make better use of the access to a small number of unconfounded samples than softmax. Once model mistakes have been detected, CBMs provide a straightforward way to improve a model via human feedback. Hence, we next evaluate the effect of interventions applied after detection of CB2M and the baselines on model performance.
In Tab. 3, we report the normalized relative improvement (NRI) on the test set to evaluate the improvement of interventions applied to the detected mistakes. Both for CUB and Parity MNIST (unbalanced), interventions can improve model performance on an instance (close to) 100%. This results in similar NRIs for all methods on the identified instances. More important, however, is the effect observed on the full dataset. Here, we can see that interventions after random selection only have a small effect. Interventions applied after the softmax baseline already cause a larger improvement, while interventions after samples detected for CB2M yield the largest improvement. Based on this, we can conclude that the mistakes detected via CB2M are indeed more suited for interventions than the baselines.
Often, intervening on a few concepts is already sufficient because they carry most of the relevant information. As human interactions are expensive, we want to avoid unnecessary human feedback and only ask for interventions on the relevant concepts. As shown in Shin et al. (2023); Chauhan et al. (2022), selecting the concepts for intervention can reduce the required human interactions by a large amount. In Fig. 3, we show the increase in performance when applying interventions after CB2M detection for a progressive number of concepts. One can observe that interventions on a few concept groups (e.g. 10) already yield a large portion of the maximum improvement (around 60%). Applying interventions beyond 19 concept groups barely shows further improvements. [^2] Performing interventions on all available concepts cannot necessarily improve class accuracy to 100%, as interventions cannot provide values for invisible concepts. Even if the presented results here are based on "full" interventions for brevity, this shows that we do not need interventions on all concepts to achieve the benefits of CB2M.
Generalization under Distribution Shift.
Lastly, we want to evaluate the benefits of CB2M when the base CBM is object to a distribution shift. To that end, we first train a CBM on Parity MNIST and then evaluate it on Parity SVHN. As seen in Tab. 4, the base model does not perform well under the shift, with a class accuracy barely over 50% (which is equal to random guessing). Nevertheless, if we add some human-generated interventions to CB2M, we can still improve the model performance under the distribution shift. However, it is notable that the improvement is smaller than in the other use cases shown in previous sections. This is most likely because model encodings on SVHN are less concentrated due to the drastic distribution shift. Still, we believe these results show the potential of CB2M other settings like online learning.
Table 4: CB2M generalization under distribution shift. The CBM is trained on Parity MNIST and evaluated on SVHN. Despite the low base model performance, CB2M can still generalize human interventions on SVHN. (Best values bold, standard deviations over 5 runs.)
Concept Acc. ($\uparrow$) Class Acc. ($\uparrow$)
Setting CBM CB2M CBM CB2M
Identified $63.1\pm 1.2$ $\mathbf{87.3}\pm 0.1$ $39.9\pm 0.3$ $\mathbf{60.8}\pm 0.4$
Full set $68.0\pm 0.9$ $\mathbf{75.3}\pm 0.4$ $51.0\pm 0.1$ $\mathbf{57.3}\pm 0.2$
4 Related Work
In the following, we highlight research related to CB2M. First, we discuss works around CBMs before highlighting some related works on model error detection.
Concept Bottleneck Models.
Concept bottleneck models as a general network architecture were popularized recently by Koh et al. Koh et al. (2020). The two staged model first generates an intermediate concept representation before generating the final task output. Since their introduction, CBMs have been extended in various ways. To reduce the dependency on fully supervised concept information, CBM-AUC Sawada & Nakamura (2022) combine explicit concept supervision with unsupervised concept learning found in self-explaining neural networks Alvarez-Melis & Jaakkola (2018). Similarly, PostHoc CBMs Yüksekgönül et al. (2022) and label-free CBMs Oikarinen et al. (2023) encompass concepts from concept libraries (e.g. with CAV Kim et al. (2018)) to require less concept supervision. Stammer et al. Stammer et al. (2022) introduces a method for learning concepts with weaker supervision based on discretizing prototype representations.
The internal concept representation is the key factor for CBM interpretability. However, a CBM sometimes encodes more than the intended concept information in its internal representation Margeloiu et al. (2021). This problem of concept leakage undermines the inherent interpretability properties, as the information provided by a concept becomes unclear Havasi et al. (2022); Mahinpei et al. (2021). As a result, GlanceNets Marconato et al. (2022) and concept embedding models Zarlenga et al. (2022) have been introduced to solve leakage problems in different ways, and CBMs have been extended to drop concept predictions when not enough knowledge is available Lockhart et al. (2022). As various CBM architectures exist, the flexibility of our presented CB2M is desirable. The only requirements to combine CB2M with other CBM architectures are access to the model encodings and the ability to apply interventions.
There has also been some research done on the intervention procedure of CBMs. By overwriting the internal concept representation of a model, humans can provide corrective feedback to the CBM. The intervention procedure introduced in Koh et al. (2020) is quite simple, selecting concepts randomly and providing interventions to all examples. Since then, Shin et al. Shin et al. (2023) presented several heuristics to order concepts for intervention, drastically reducing the number of concepts necessary to correct the model. Similarly, SIUL Sheth et al. (2022) uses Monte Carlo Dropout to estimate the importance of concepts and, therefore, the need for interventions. Interactive CBMs Chauhan et al. (2022) extend this idea even further by providing a policy to optimize concept selection over several examples under consideration of intervention costs Chauhan et al. (2022). Still, these works only consider the ordering of concepts for interventions, whereas we explicitly consider the example selection. And even more importantly, neither of these methods changes the fact that vanilla CBM interventions only have a one-time effect. They can, however, be combined with CB2M to improve the intervention procedure even more.
Uncertainty Estimation for Error Detection.
One use case of CB2M is to detect potential model mistakes (which can then be improved via interventions). Detecting data points where the model performs poorly is often touched upon under the topic of uncertainty estimation. While the construction of uncertainty-aware networks can often provide benefits in terms of mistake detection Gawlikowski et al. (2021), our work is more related to methods that do not make particular assumptions about the model architecture. This is essential to ensure that our method can be combined with different CBM architectures.
One popular approach to detect model mistakes is using the softmax probabilities of the model Hendrycks & Gimpel (2017). One can predict if a model makes a mistake by setting a threshold to the softmax probability of the most likely class. As a different example, TrustScore Jiang et al. (2018) measures the agreement between the model output and a KNN classifier for the same task to detect model mistakes. Even if these models are also used to detect model mistakes, they are not specifically tailored to CBMs. As a CBM consists of two different networks, errors can occur in either. Both the softmax baseline and TrustScore only look at the model prediction as a whole, while our method is able to specifically detect mistakes related to the bottleneck network (which can then be corrected via interventions). In contrast, neighborhood uncertainty classifiers Ramalho & Miranda (2019) learn a neural network on top of a KNN of latent model representations to predict uncertainty. This method also uses a latent representation of the model. However, we do not learn a neural network on top of similarity information, keeping our technique simpler and more flexible. That way, it can even be adapted if more details about model mistakes arrive at model deployment, which happens every time a human provides an intervention.
5 Conclusion
In this work, we have introduced CB2M, a flexible extension to CBM models. We have shown that the two-fold memory of CB2M can be used to generalize interventions to previously unseen datapoints, thereby overcoming the issue of current one-time intervention approaches without the necessity of further human interactions. Additionally, we have demonstrated that CB2M can be utilized to detect model mistakes prior to any human interactions, allowing query humans to efficiently provide interventional feedback in a targeted manner. CB2M can be combined with any concept model architecture to greatly improve interventions and make human interactions more efficient.
There remains room to improve CB2M in the future. First, one could make the memory of CB2M differentiable Plötz & Roth (2018). This would allow to learn parameters like $t_{d}$ directly instead of relying on heuristics. Therefore, a differentiable memory could enhance the coverage of intervention generalization. Here, we present CB2M with generalizing interventions only from the closest mistake, i.e., using $k=1$. To further boost the performance gains with CB2M, it is also conceivable to combine interventions from multiple mistakes. A simple way to achieve this could be, e.g., weighted averaging, but more complex aggregation methods are also possible. Such aggregation techniques would furthermore allow condensing information in the memory into prototypes, keeping the memory smaller and better understandable. As shown, CB2M excel particularly when the model makes systematic errors, like with confounded data or when the model fails to learn a specific class. However, if the model performance is really good and the only mistakes are a few singular outliers (potentially close to the decision boundary), CB2M do not work as well. If there are no similar model mistakes, detection solely based on mistakes in the memory is not sufficient. Possible directions to apply CB2M in such circumstances could be integrating correctly classified instances to the memory, allowing for a better inlier detection. As it is crucial for all work on interactive machine learning, we encourage future research in the context of malicious human interactions, specifically how to prevent them from misusing the memory of CB2M. Finally, an interesting future direction is the combination of CB2M with other concept-based models, for example CEM Zarlenga et al. (2022), post-hoc CBMs Yüksekgönül et al. (2022) or even tabular CBMs Zarlenga et al. (2023).
Acknowledgments This work benefited from the Hessian Ministry of Science and the Arts (HMWK) projects ”The Third Wave of Artificial Intelligence - 3AI”, "The Adaptive Mind" and Hessian.AI, the Hessian research priority program LOEWE within the project WhiteBox, the "ML2MT" project from the Volkswagen Stiftung as well as from the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215).
Appendix A Appendix
A.1 Additional Training Details.
On the Parity MNIST, SVHN, and C-MNIST datasets, we used an MLP both for the bottleneck and predictor networks. The bottleneck is a two-layer MLP with a hidden dimension of 120 and ReLU activation functions, while the predictor is a single-layer MLP. The bottlenecks are trained using the specific dataset’s respective training and validation sets. Notably, for the Parity MNIST (unbalanced), the training unbalance is not present in the validation data. For the generalization and mistake detection experiments on C-MNIST, the human-provided interventions are from the unconfounded data, which is 10% of the original C-MNIST test dataset, which was neither used for training nor evaluation. Evaluation is done on the remainder of the test set.
A.2 Algorithm for Intervention Generalization.
For reference, we present the algorithm for intervention generalization (Alg. 2). To check whether a newly obtained test instance $\hat{x}$ should be automatically updated with an intervention from memory, we do the following: First, the mistake $x^{\prime}$ with the smallest encoding distance to the new datapoint is retrieved from the mistake memory $M^{m}$. Then, if the distance between encodings of both samples is smaller than $t_{d}$, we check if there exists an associated intervention $i$ to $x^{\prime}$, i.e., if there is an $i\in\mathcal{M}^{i}$ with $\alpha(x_{e}^{\prime},i)$. If so, we apply this intervention. Otherwise, we do not modify $\hat{x}$. Afterward, the predictor network of the CBM performs its prediction with the potentially modified $\hat{x}$ as usual.
1:  $\hat{x}\leftarrow\texttt{New unseen instance}$
2:  Obtain $\hat{x}_{e}$ through $g$
3:  find $x^{\prime}\in\mathcal{M}^{m}$ with minimal $d(\hat{x}_{e},x^{\prime}_{e})$
4:  if $d(\hat{x}_{e},x^{\prime}_{e})<t_{d}$ then
5:     if $\exists i\in\mathcal{M}^{i}:\alpha(x_{e}^{\prime},i)$ then
6:        $x\leftarrow x|i$
7:     end if
8:  end if
9:  Model Output: $y=f(x)$
Algorithm 2 Generalize Interventions to Unseen Images Given: CBM with bottleneck $g$ and predictor $f$, threshold parameter $t_{d}$ and a memory $\mathcal{M}=(\mathcal{M}^{m},\mathcal{M}^{i})$ of reference mistakes with respective interventions.
A.3 Results on Parity MNIST
For reference, we provide results when applying CB2M to Parity MNIST. The performance of the base CBM on this task is already pretty good, as it achieves a concept accuracy of $98.9$% and a class accuracy of $97.7$%. The few errors that the model makes are due to singular outliers. As discussed in Sec. 5, the CB2M performs well when the model is subject to some kind of systematic error, e.g., when the model is subject to a shift in data distribution or due to data imbalance at training time. When model mistakes are just a few individual examples, which are getting confused with different classes, CB2M does not perform as well (Tab. 6, 6). As the base CBM performance is already good, further intervention generalization is not suitable, as the remaining model mistakes are not similar to each other (Tab. 7). Further adjustments like including positive examples in the memory or using an explicit view on mistake density could potentially improve results in these situations.
Table 5: Detection of model mistakes on Parity MNIST. For mistake detection on models with a low error rate (with errors being outliers close to the decision boundaries), CB2M performs worse than softmax. (Best values bold, standard deviations over 5 runs.)
Table 6: Interventions after detection on Parity MNIST. NRI on identified instances and full set. Interventions successfully improve identified instances. However, worse detection than softmax results in smaller improvement via CB2M. (Best values bold, standard deviations over 5 runs.)
Random Softmax CB2M
AUROC ($\uparrow$) $49.0\pm 0.4$ $\mathbf{93.3}\pm 0.2$ $64.6\pm 1.0$
AUPR ($\uparrow$) $97.4\pm 0.0$ $\mathbf{99.8}\pm 0.0$ $98.8\pm 0.1$
Setting Random Softmax CB2M
Identified $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$
Full Set $1.6\pm 0.7$ $\mathbf{57.6}\pm 1.5$ $5.9\pm 2.9$
Table 6: Interventions after detection on Parity MNIST. NRI on identified instances and full set. Interventions successfully improve identified instances. However, worse detection than softmax results in smaller improvement via CB2M. (Best values bold, standard deviations over 5 runs.)
Table 7: Generalization of CB2M does not impact results on Parity MNIST. As model mistakes are not similar to each other, no instances have been identified for intervention generalization, therefore applying CB2M does not impact model performance. (Best values bold, standard deviations over 5 runs.)
Setting CBM CB2M
Concept Acc. ($\uparrow$) Identified $\emptyset$ $\emptyset$
Full set $\mathbf{98.9}\pm 0.0$ $\mathbf{98.9}\pm 0.0$
Class Acc. ($\uparrow$) Identified $\emptyset$ $\emptyset$
Full set $\mathbf{97.7}\pm 0.0$ $\mathbf{97.7}\pm 0.0$
A.4 Further Results on Parity MNIST (unbalanced).
The unbalanced version of Parity MNIST is generated by dropping 95% of the training data of one class. In the main paper, we exemplarily showed the results when removing digit 9. In Tab. 8, we show the average results for all other digits. The base mode does not capture the training imbalance properly in three cases, resulting in larger standard deviations for all results.
Table 8: Further results on Partiy MNIST (unbalanced). Results of all main experiments for all versions of the Parity MNIST (unbalanced) dataset (where the digits 0 to 8 where the underrepresented digits respectively). (Average and standard deviation over unbalance with digits 0 to 8.)
Mistake Detection
Random Softmax CB2M
AUROC ($\uparrow$) $49.5\pm 1.0$ $\mathbf{91.2}\pm 7.2$ $83.8\pm 10.77$
AUPR ($\uparrow$) $92.8\pm 3.5$ $\mathbf{99.3}\pm 0.4$ $98.9\pm 0.3$
Performance after Interventions (NRI)
Setting Random Softmax CB2M
Identified $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$
Full Set $24.3\pm 21.3$ $70.7\pm 13.8$ $\mathbf{75.6}\pm 14.6$
Generalization of Interventions
Setting CBM CB2M
Concept Acc. ($\uparrow$) Identified $91.8\pm 2.1$ $\mathbf{97.4}\pm 2.3$
Full Set $98.4\pm 0.0$ $\mathbf{98.6}\pm 0.3$
Class Acc. ($\uparrow$) Identified $37.2\pm 26.4$ $\mathbf{87.1}\pm 11.4$
Full Set $92.9\pm 3.5$ $\mathbf{95.0}\pm 2.9$
A.5 Further Details on Generalization Results
In Sec. 3.2, we show the generalization capabilities of CB2M on various datasets. To further detail these results, the number of generalized interventions is presented in Tab. 9. This describes to how many unseen examples the human interventions have been generalized. The standard deviation is generally relatively large, especially for the CUB dataset. This is most likely caused by the selection of the threshold parameter $t_{d}$. The number of generalized interventions for the parity MNIST to SVHN dataset is considerably larger, as this dataset has more datapoints, and the model makes more mistakes after the distribution shift.
Table 9: Number of generalized interventions for the different datasets. For SVHN, the number of model mistakes is considerably larger, therefore there are more possible generalizations. (Average and standard deviations over 5 runs.)
Dataset Number of Intervention Generalizations
CUB $289.4\pm 215.5$
Parity MNIST (unbalanced) $416.2\pm 206.5$
Parity C-MNIST $913.4\pm 342.8$
Parity MNIST to SVHN $7809\pm 512$